A solution contains 4.9 grams of sugar dissolved in 23.6 grams of water. What is the mass % of the solution?


1
Expert's answer
2021-04-01T06:03:55-0400

.w(sugar)=m(sugar)/M(solution)=4.9/28.5=0.101or10.1